OpenStudy (anonymous):

z1=1+sqrt3 i part 1: Find the four angles that define the fourth roots of the number z1. part 2: What are the fourth roots of z1= sqrt3 +1?

OpenStudy (psymon):

Well, we need to transform this a+bi form into: \[r(\cos \theta + isin \theta)\]Now hopefully you have conversions for these somewhere where you'd be able to do them on your own, but these are the two conversions between rectangular and polar that we'll use: \[r = \sqrt{a ^{2}+b ^{2}}\]which you may recognize as pythagorean theorem/distance formula. The second is: \[\tan \theta = \frac{ b }{ a }\] So with those conversions, you think you'd be able to start off and get r for me?

OpenStudy (anonymous):

r=2

OpenStudy (psymon):

Awesome. So now \[\tan \theta = \frac{ \sqrt{3} }{ 1 }\]Do you know what angle for tangent gets sqrt(3)?

OpenStudy (anonymous):

60 degrees

OpenStudy (psymon):

Sounds right to me. So now we have this: \[2(\cos60+isin60)\] Now this is the polar form, but we need roots. So this is how we find the roots. \[\sqrt[n]{r}[\cos (\frac{ \theta }{ n }+\frac{ 2k \pi }{ n })+isin(\frac{ \theta }{ n }+\frac{ 2k \pi }{ n })]\] The k starts from 0 and goes up to n-1 where n is the root we want. So if we want 4th roots, we want to START from k = 0 and n = 4. So this is our first 4th root: \[\sqrt[4]{2}(\cos15+isin15)\] To get the next 3 roots, make k = 1, 2, 3 and get your new angles :3

OpenStudy (anonymous):

:/ wow thats confusing

OpenStudy (psymon):

Since you're dealing with degrees instead of radians, probably best to write it like: \[\frac{ 360k }{ n }\]

OpenStudy (psymon):

Okay, do you understand how we got thefirst answer, with 4th root 2 and 15 degrees?

OpenStudy (anonymous):

not really

OpenStudy (psymon):

Ah. Okay, well how about the way we got 2(cos60+isin60)?

OpenStudy (anonymous):

yeah I understand that it's just this whole roots thing thats super confusing

OpenStudy (psymon):

Right. I tried to put it in formula form, but that's prob a bit much xD Alright, so whatever root we want, we take r to that root and divide the angles by the number root. So since we want 4th root, I took the 4th root of 2 and divided the angles by 4.

OpenStudy (anonymous):

okay so for the next root angle thing, would I just divide 15 by 4?

OpenStudy (psymon):

Nope. the 4th root 2 and the 15 degree one is our first root. The next roots are where the whole 360k/n part comes in. Since we have a 4th root, n = 4 and that 360 thing becomes 90k. So basically, the next 3 roots we want are found by adding 90 degrees. So if we want the 2nd root we add 90 to get 105. The next root we add 90 again to get 195. The final root we add 90 again to get 285. Those are your 4 roots :3 the 4th root(2) stays the same.

OpenStudy (anonymous):

I get what you are saying, but I'm not sure what numbers to punch into my calculator to get the next root.

OpenStudy (psymon):

You don't punch numbers in, just leave it all in degrees or radical form. 4th(2)(cos15+isin15) 4th(2)(cos105+isin105) 4th(2)(cos195+isin195) 4th(2)(cos285+isin285) Those would be your answers.

OpenStudy (anonymous):

ohhhh! Okay! Thanks! So for the next part I have r=2 and the modulus is 30 degrees, so what next?

OpenStudy (psymon):

What does it want you to do with it?

OpenStudy (anonymous):

find the fourth roots of z1=sqrt3+i

OpenStudy (psymon):

Okay, cool, r = 2 modulus is 30. So I think it's best to keep it in radians, so pi/6. Up toyou, unless it forces you to use degrees xD But yeah, same process, we take the 4th root of r and divide the angles by 4.

OpenStudy (anonymous):

okay so the first root is 4root2(cos(7.5)+isin(7.5)) right?

OpenStudy (psymon):

correct. Now we take 360 and divide it by the root number, which gives us 90 degrees again. So now keep adding90 degrees until you have all 4 :P

OpenStudy (anonymous):

so the next roots will be 97.5, 187.5, and 277.5, right?

OpenStudy (psymon):

That's all there is to it.

OpenStudy (anonymous):

yay! Thanks again! You're my life saver!

OpenStudy (psymon):

As long as you understand it then its all good xD